OpenStudy (thebadyboy):

Solve the equation. x + 1/6 = 6 A. x = –6 1/6 B. x = –5 5/6 C. x = 5 5/6 D. x = 6 1/6

OpenStudy (thebadyboy):

1/6 etc. are fractions

OpenStudy (rhystic):

bring the 1/6 to the other side so you have x = 6 - 1/6 which is the same as 5 5/6

OpenStudy (kendricklamar2014):

1. Subtract 1/6 from both sides @Thebadyboy

OpenStudy (thebadyboy):

ok and @KendrickLamar2014 its me the user822 i made new account

OpenStudy (kendricklamar2014):

:)

OpenStudy (kendricklamar2014):

So the equation will look like this:\[x + 1/6 - 1/6 = 6 - 1/6\]

OpenStudy (thebadyboy):

ok so a

OpenStudy (thebadyboy):

right?

OpenStudy (kendricklamar2014):

I was thinking D

OpenStudy (thebadyboy):

ok u seem smarter

OpenStudy (thebadyboy):

hey follow me

OpenStudy (kendricklamar2014):

I did, and I wrote a Testimonial :)

OpenStudy (thebadyboy):

thx

OpenStudy (kendricklamar2014):

\[No Problem\]